OpenStudy (anonymous):

Why does the addition sign change in 9^-(n+2) when you convert it to 1/9^(n-2)?

OpenStudy (campbell_st):

there is now change \[9^{-(n +2)} = \frac{1}{9^{n + 2}}\] for a proof \[\frac{1}{9^n} \times \frac{1}{9^2} = \frac{1}{9^{(n + 2)}} = 9^{-(n + 2)}\]

OpenStudy (anonymous):

so what is happening in this step? http://tutorial.math.lamar.edu/Classes/CalcII/Series_Special_files/eq0020MP.gif

OpenStudy (anonymous):

oh. wait i think i get it. nvm